Here is a simplified version of the problem I am working on 1*10^-14=(x)(y)
considering, x is 1000 times greater than y. I tried to set up the work like this: 1*10^-14=(1000x)(x), so 1*10^-14=1000x^2. I was easily able to solve for x, but (x*1000)(x) does not equal 1*10^-14. Can you help me?

Answer by stanbon(75887)

1*10^-14=(x)(y) considering, x is 1000 times greater than y.
--------
I tried to set up the work like this: 1*10^-14=(1000x)(x), so 1*10^-14=1000x^2. I was easily able to solve for x
-----
10^-14 = 1000x^2
x^2 = 10^-17
x = 10^-8.5
Solve for "y":
y = 1000x
y = 1000*10^-8.5
y = 10^-(5.5)
================
Cheers,
Stan H.
================
----------------------------------
Cheers,
Stan H.